OpenStudy (anonymous):

Integral of x^3/(radical(1+x^4)

OpenStudy (anonymous):

u=1+x^4 du=4x^3 du/4=x^3?

OpenStudy (anonymous):

Very close: \[ u=1+x^4\\ du=4x^3dx\\ \frac {du}4=x^3dx \]

OpenStudy (anonymous):

so is it du/u?

OpenStudy (anonymous):

du/4u*

OpenStudy (anonymous):

\[ \sqrt{1+x^4}\to \sqrt{u} \]

OpenStudy (anonymous):

right sorry. so du/4 is acceptable?

OpenStudy (anonymous):

Yes,

OpenStudy (anonymous):

so would it be 1du/4 radical(u) then you integrate it?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

um so,4u^1/2/2?